OpenStudy (anonymous):

Two positive numbers have a difference of 8. The larger number is 1 more than twice the smaller. Find the two numbers

OpenStudy (theviper):

Let the first no be \(x\) then according to the question other no. will be 2x+1 Then 2x+1-x=8 x+1=8 x=8-1=7 & 2x+1=2(7)+1=14+1=15 Therefore the no. are 7 & 15.

OpenStudy (theviper):

Got it??

OpenStudy (anonymous):

i wasnt getting the second part thank you

OpenStudy (theviper):

:)